# Investigating Bias in LLM-Based Bias Detection: Disparities between LLMs and Human Perception

### v1

@@ Appendix B Latent Topic Construction
Latent Topic Construction
Inspired by IndiVec Lin et al. (2024b), we follow them to prompt ChatGPT to construct a fine-grained media bias indicators dataset based on the dataset Flipbias. The indicators are the summarised key points that may reflect media bias from each article. To better arrange the topics covered in these articles, we do the strict clustering through Hierarchical Clustering based on Euclidean distance on the indicators that got from the dataset Flipbias. We used AgglomerativeClustering in the package Scikit Learn, and the distance threshold is set to 2. The embedding of the indicators is obtained from OpenAIEmbeddings. Finally, 19671 indicators are clustered into 152 clusters. Each cluster stands for a latent topic.
@@
Indicators
Interpreted Topic
"Provides figures and quotes from individuals involved in the issue", "The article cites statements from various individuals involved in the case, including lawyers, politicians, and advocacy groups.", "The text quotes various experts and government officials to support its claims.", "cites tweets and quotes from Trump, experts, and state officials to support the claims made", "Quotes from various food experts and diplomats.", "The article cites multiple sources, including government documents and quotes from officials." … …
Comprehensive Use of Quotes and Citations in Journalism
"Describes President Obama’s decision as "benighted" and "cowardly" while praising President Trump’s decision", ""swipes at Joe Biden," "knocks primary rival Bernie Sanders," "gripes about former President Barack Obama"", "frames the issue as a result of understaffing and mismanagement, blames the Obama administration, and highlights the need to protect the president", "Celebratory tone towards Obama, sarcastic and mocking tone towards Democrats", "Portrays Democrats as wanting a grander celebration, mocks Obama and the holiday", "Frames the decision as a potential unwinding of an Obama executive action, includes criticism from Democrats and environmental groups", "Describes the tough-on-crime approach as a reversal of Obama’s "Smart on Crime" policy, implying a negative change" … …
Diverse Perspectives on President Obama’s Policies and Actions
"Mentions celebrations and security measures in various cities", "Mentions specific incidents of terrorism and security measures in different cities", "Presents the incident as a terrorist attack and highlights the victims’ nationalities", "Mentions previous vehicle attacks and quotes from witnesses", "Provides examples of other major music event attacks", "Mentions the arsenal of weapons and ammunition recovered, suggesting the possibility of an accomplice", "Mentions the London subway station fire as a terrorist incident caused by an improvised explosive device", "The article provides examples of previous attacks and the use of improvised explosive devices." … …
Analysis of Recent Terrorist Attacks and Security Measures in Various Cities
Table 6: Clustered Indicators and Interpreted Topics.
Please refer to Table 6.
@@
Interpretation of Top and Bottom 5 latent topics (ranked by BTI-2 values)
BTI-1 BTI-2 Frequency
Trump’s Clashes with Federal Law Enforcement and Media
-0.17 0.11 80
Examining Controversial Tactics: Dissecting Allegations and Defenses in Recent Political Affairs
0.18 0.10 64
Analysis of Congressional Dynamics: Trump’s Strategy, Witness Battles, and Financial Focus
-0.05 0.09 72
Bipartisan Cooperation in Senate: Struggles and Progress
-0.05 0.08 68
Unveiling the Constitutional Crisis: Examining Government Overreach and the Erosion of Rights
0.25 0.07 72
Understanding Textual Analysis: The Importance of Examples and Analogies
0.18 -0.05 59
Statewide Controversies: Voter Rights, Criminal Justice, and Transition Integrity
0.18 -0.06 60
Analyzing Political Discourse: Insights from Trump Administration and Beyond
-0.02 -0.08 76
Examining Biased Reporting in Political Discourse: Imbalance in State of the Union Addresses
0.20 -0.09 96
Critical Discourse Analysis of Media Portrayal on Trump’s Governance
-0.01 -0.10 91
Table 7: Interpretation of Top and Bottom 5 Latent Topics on FlipBias.
Please refer to Table 7.
@@ C.1 Details of Prompts to Isolate Bias
We randomly selected 4 Left-Center-Right triples from the dataset Filpbias, and then used the titles as the instances of few-shot instruction, which are listed in Table 8.
Text
Label
Trump Accuses His Justice Department, FBI Of Favoring Democrats
Left
Explosive memo released as Trump escalates fight over Russia probe
Center
Trump accuses FBI, DOJ leadership of bias against Republicans and in favor of Dems
Right
Shutdown truce just delays Trump’s big dilemma
Left
Winners and losers from the government shutdown
Center
Centrists break Senate logjam, pave new path for ‘common sense’ bipartisanship
Right
North Korean insults to U.S. leaders are nothing new — but Trump’s deeply personal reactions are
Left
Trump trades ’short and fat’ barb with N Korea’s Kim
Center
Trump Take To Social Media To Hit Back At ’Short and Fat’ Kim Jong-un
Right
After 16 Futile Years, Congress Will Try Again to Legalize ‘Dreamers’
Left
The clock is ticking’: Graham and Durbin urge action on bipartisan DREAM Act by the end of September
Center
Republican Sen. Cory Gardner agrees to support bipartisan Dream Act after Trump rescinds DACA
Right

### v2

@@ Appendix B Latent Topic Construction
Inspired by IndiVec Lin et al. (2024), we prompted ChatGPT to construct fine-grained media bias indicators using the Flipbias dataset. These indicators summarize key points that may reflect media bias in each article. To organize the topics covered in these articles more effectively, we performed strict clustering through Hierarchical Clustering based on Euclidean distance applied to the indicators extracted from Flipbias. We utilized AgglomerativeClustering from the Scikit Learn package, setting the distance threshold to 2. The embeddings of the indicators were derived from OpenAIEmbeddings. Ultimately, 19,671 indicators were clustered into 152 clusters, each representing a latent topic.
@@
Details of the clustered indicators are provided in Table 6.
@@
Rankings of latent topic cases based on BTI-2 values are shown in Table 7.
@@ C.1 Details of Prompts to Isolate Bias
Table 8: Examples of Article Continuation
In Section 5.1, we discussed the methods to debias LLMs. Here we provie details of these debaising methods.
Bias Label Explanation
In Bias Label Explanation (BLE) method, we adopt explanations as listed in Table 8.
@@
In Few-shot Instruction method, we randomly selected 4 Left-Center-Right triples from the dataset Flipbias and then used the titles as the instances of few-shot instruction, which are listed in Table 9.
Text Label
Trump Accuses His Justice Department, FBI Of Favoring Democrats Left
Explosive memo released as Trump escalates fight over Russia probe Center
Trump accuses FBI, DOJ leadership of bias against Republicans and in favor of Dems Right
Shutdown truce just delays Trump’s big dilemma Left
Winners and losers from the government shutdown Center
Centrists break Senate logjam, pave new path for ‘common sense’ bipartisanship Right
North Korean insults to U.S. leaders are nothing new — but Trump’s deeply personal reactions are Left
Trump trades ’short and fat’ barb with N Korea’s Kim Center
Trump Take To Social Media To Hit Back At ’Short and Fat’ Kim Jong-un Right
After 16 Futile Years, Congress Will Try Again to Legalize ‘Dreamers’ Left
The clock is ticking’: Graham and Durbin urge action on bipartisan DREAM Act by the end of September Center
Republican Sen. Cory Gardner agrees to support bipartisan Dream Act after Trump rescinds DACA Right
@@ Appendix D Debiasing Results
Appendix E More Discussions
E.1 Distinguish from Related Works
Existing research has explored political bias in LLMs. Here, we differentiate our contributions from key related works:
Taubenfeld et al. (2024) examines LLMs in simulating political debates, revealing conformity to inherent social biases despite specific directions. Taubenfeld et al. (2024)focuses on interaction simulation, whereas our research centers on bias detection, emphasizing end-to-end and fine-grained analyses. Rozado (2024) analyzes bias through 11 political orientation tests, while our study highlights limitations of orientation tests and provides robust quantitative analysis based on extensive datasets, offering a broader perspective on LLM biases. Urman and Makhortykh (2023) investigates LLM-Chat Models’ responses to predefined queries, focusing on non-responses and false responses. While related to the political domain, it primarily addresses jailbreaking and harmful effects. Our research questions are more specific, targeting systematic bias detection. Motoki et al. (2024) evaluates ChatGPT’s responses to ideological questions. It focuses solely on ChatGPT, whereas our work encompasses a broader range of LLMs and addresses comprehensive research questions, providing a more extensive analysis.
These studies contribute to understanding political bias in LLMs. However, our work stands out by offering a more systematic exploration, addressing four comprehensive research questions, employing intricate experimental designs, and analyzing a broader range of LLMs, thus significantly extending the current body of research.
E.2 Exploration of Different Embeddings
In Section 3.2, we explored the embedding-based similarity matching method using embeddings from the GPT-3.5 model. Here, we extend our investigation to include another embedding source: sentence-t5-base [^3] https://huggingface.co/sentence-transformers/sentence-t5-base (T5-Base). The continuation results using T5-Base embeddings are summarized in Table 12. The calculation of left and right percentages in the table follows the methodology detailed in Figure 4.
From Table 12, we observe similar trends across different prefix lengths as shown in Fig. 4, although there are slight variations in predictions for prefix length = 320. Overall, the findings indicate a predominant left-leaning trend in continued articles, consistent with our earlier observations using GPT-3.5 embeddings.
E.3 Result of Article Continuation based on classifier
20
(left/right)
40
(left/right)
80
(left/right)
160
(left/right)
320
(left/right)
ZERO-SHOT 41.8/58.2 19.3/80.7 20.0/80.0 41.1/58.9 27.3/72.7
FEW-SHOT 70.5/29.5 53.6/46.4 53.3/46.7 51.3/48.7 48.0/52.0
Table 10: Results of article continuation based on classifier (GPT 3.5).
In this experiment using gpt-3.5-turbo-0613 as a classifier, we have two settings. In the setting zero-shot, we directly prompt ChatGPT3.5 to answer whether the given continuation is Left, Center, or Right, in the setting few-shot setting, we give the same reference as we compared in embedding-based similarity matching, which is a triple of Left-Center-Right articles that are of the same event as our continuation. The results of the ratio Left/Right are shown in Table 10.
E.4 Article Continuation Examples
Prefix-20 Example:
Prefix: WASHINGTON (Reuters) - President Donald Trump and his Republican allies in Congress on Friday escalated a campaign against
Suffix: Special Counsel Robert Mueller and the ongoing Russia probe, as an explosive memo was released that alleges misconduct within the FBI and Justice Department. The memo, written by Republican lawmakers, accuses the FBI of abusing its surveillance powers in obtaining a warrant to spy on a former Trump campaign aide…
Prefix-40 Example:
Prefix: WASHINGTON (Reuters) - President Donald Trump and his Republican allies in Congress on Friday escalated a campaign against U.S. law enforcement agencies over their probe of Trump’s ties to Russia, releasing a disputed memo that
Suffix: accuses the FBI of bias and misconduct. The explosive memo, which was released by the House Intelligence Committee, has sparked intense controversy and further deepened the divide between the Trump administration and the intelligence community. …
Prefix-80 Example:
Prefix: … Trump’s ties to Russia, releasing a disputed memo that the FBI warned was misleading and inaccurate. The previously classified document, written by Republicans on the House of Representatives intelligence committee, alleges that the federal probe of potential collusion between Trump’s 2016 presidential campaign and
Suffix: Russian officials was politically motivated and based on questionable evidence. The release of the memo marks a significant turning point in the ongoing battle between Trump and the intelligence community. Trump has long criticized the investigation into Russian interference in the 2016 election, calling it a "witch hunt" and claiming that it is a politically motivated attempt to undermine his presidency. …
Table 11: Examples of Article Continuation
20
(left/right)
40
(left/right)
80
(left/right)
160
(left/right)
320
(left/right)
GPT3.5 52.8/47.2 53.1/46.8 54.4/45.6 54.8/45.2 41.3/58.7
T5-base 50.0/50.0 49.9/50.1 52.8/47.2 51.7/48.3 57.6/42.4
Table 12: Comparison of Two Embeddings (GPT3.5 v.s. T5-Base) Results in Article Continuation Experiments
In Section 3.2, we adopt GPT-3.5 to conduct article continuation. We first report the average suffix length for each setting as follows: 490.1, 487.5, 479.0, 463.7, and 473.9 for prefixes with lengths of 20, 40, …, 320, respectively. Due to the strong capability of GPT-3.5, the generated suffixes are quite consistent with the prefix. Table 11 shows the randomly selected examples in different prefix settings of article continuations.
E.5 Finetuning Other LLMs
In addition to the finetuning debiasing results of ChatGPT 3.5 reported in Table 4, we examined the finetuning debiasing method on a smaller LLM, specifically LLaMa-2-7B-Chat.
We report the BTI-1 and BTI-2 scores for LLaMa2 in Table 13, where:
  • LLaMa2: LLaMa-2-7B-Chat without finetuning.
  • LLaMa2 LCR-FT: LLaMa-2-7B-Chat finetuned according to the setting described in Section 5.2, with 300 articles evenly distributed among left-label, center-label, and right-label categories (LCR-FT).
  • LLaMa2 Finetune (Right Leaning data): LLaMa-2-7B-Chat finetuned with 300 right-leaning data, where grounded center articles are labeled as left, and grounded right articles are labeled as center.
(a) LlaMa2
(b) LlaMa2 LCR-FT
Figure 12: BTI distribution of Topics on FlipBias dataset for LlaMa2 and LlaMa2 LCR-FT.
We further report the BTI distribution of LLaMa2 and LLaMa2 LCR-FT in Fig. 12. We observe that although the averaged BTI-1 scores do not exhibit significant changes in Table 13 before and after finetuning, upon examining the topic-level distribution (refer to Fig. 12), we notice a more centralized BTI-1 distribution.
Model BTI-1
LLaMa2 0.04
LLaMa2 LCR-FT -0.024
LLaMa2 Finetune (Right Leaning data) 0.02
Table 13: BTI-1 of LLaMa2 and Finetuning Methods